OpenStudy (anonymous):

There are no real solutions for a quadratic equation when the discriminant is: a. zero b. negative c. positive d. less than five e. greater than eight

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

@UnkleRhaukus what does this mean?

OpenStudy (unklerhaukus):

i have draw a quadratic equation (a parabola) with no real solutions- (it dosen't cut the x-axis)

OpenStudy (anonymous):

is the answer b ?

OpenStudy (unklerhaukus):

consider the quadratic formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] the quadratic formula is used to find the solution(s) to \[ax^2+bx+c=0\] the solutions are y=0 when x= a solution the discriminate is \(\Delta={b^2-4ac}\) what are three options for the discriminate that correspond to different types of solution

OpenStudy (anonymous):

i have no idea

OpenStudy (unklerhaukus):

\[\left\{\begin{array} \\\Delta>0\\ \Delta=0\\ \Delta<0\end{array}\right.\], see what happens if to the quadratic formula in each of these cases